# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ 6 Numerical Experiments
In this section, we conducted numerical experiments to examine the performance of our proposed algorithms. All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package [Besancon2022-lx] [^2] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
We compare BregFW and BregAFW, which use the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, and MD (their notation is described in Appendix E). Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for the away-step FW algorithm with the open loop. Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see [Braun2022-ew]. We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, $\gamma_{\max}=1$, and the termination criterion $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle\leq 10^{-7}$ throughout all experiments.
@@
Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
@@ 6.1 $\ell_{p}$ Loss Problem
We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Let $f(x)=\|Ax-b\|_{p}^{p}$. The gradient $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi:=f$.
We consider real-world data using gas sensor data.§§§The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$. We use the $\ell_{2}$ constraint $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|^{2}\leq b_{\max}\}$. We compared BregFW with EucFW and OpenFW because ShortFW does not converge. The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000. The initial point $x_{0}$ is an all-ones vector. We set $p=1.1$, i.e., $f$ is $\mathcal{C}^{1}$ not $\mathcal{C}^{2}$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $b_{\max}=130$ and $b_{\max}=200$, respectively. Because $\nabla f$ is not Lipschitz continuous on $P$ from lacking $\mathcal{C}^{2}$, ShortFW and EucFW do not converge. The primal gap and the FW gap by BregFW are the smallest among these step-size strategies because only BregFW has a theoretical guarantee for this problem.
@@ 6.2 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic]. To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$, where $P\subset\mathbb{R}^{n}$ is a compact convex set. Let $\phi$ be defined by $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1]. In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. Therefore, $L\geq\rho/\sigma$ holds with $\sigma=1$.
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized it to have norm 1 (random seed 1234). We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $(m,n)=(1000,10000)$ and $K=2000$ and for $(m,n)=(2000,10000)$ and $K=2000$, respectively. The primal gap and the FW gap by the adaptive Bregman step-size strategy are the smallest among these step-size strategies. In both cases, BregFW stopped before the 1000th iteration.
@@ 7 Conclusion and Future Work
We propose FW algorithms with an adaptive Bregman step-size strategy for a smooth adaptable function, which does not require Lipschitz continuity of its gradient. We have established sublinear to linear convergence rates depending on various assumptions for convex and nonconvex optimization. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes. For the away-step FW algorithm, we have also established global convergence for convex optimization under the HEB condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments show that our algorithms outperform existing algorithms.
Limitation and future work:
@@
Our step-size strategy for a smooth adaptable function requires the value of $\nu$, which can be estimated, because we use the Bregman distance. Theoretically, linear convergence of the FW algorithm always holds for convex optimization if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$. For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, sublinear convergence. The condition $q=1+\nu$, which is derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$. On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, which is the special case of the HEB condition with $q=2$. Since this is derived from the weak convexity of $f$, relaxing the quadratic growth condition may not be possible. In that regard, one direction is to consider an extension using the difference of convex functions (DC) optimization. Recently, Maskan et al. [Maskan2025-oa] incorporated DC optimization into FW algorithms. While they established sublinear convergence, our algorithms may exploit DC optimization techniques.
Acknowledgement
@@
This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903. Research reported in this paper was also partially supported the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
Appendix A Results from Smooth Adaptable Property
A.1 Properties of Bregman Distances
@@
Example A.1.
Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., [Bauschke2017-hg, Example 1], [Bauschke1997-vk, Section 6], and [Dhillon2008-zz, Table 2.1].
@@
    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $00\log 0=$ 0 and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler divergence [Kullback1951-yv].
@@
    The Burg entropy $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}_{++}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence [Itakura1968-en].
@@
    Let $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. The Bregman distance $D_{\phi}$ is used in phase retrieval [Bolte2018-zt], low-rank minimization [Dragomir2021-rv], NMF [Mukkamala2019-mk], and blind deconvolution [Takahashi2023-uh].
We recall the triangle scaling property for Bregman distances from Hanzely et al. [Hanzely2021-sz, Section 2], where several properties of the triangle scaling property are also shown.
Definition A.2 (Triangle Scaling Property [Hanzely2021-sz, Definition 2]).
@@
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (A.1)
Now, substituting $z\leftarrow x$ on the left-hand side of (A.1), we obtain
@@
We will now show that a weaker version of the triangle scaling property can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear.
Lemma A.3.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, there exists $\nu>0$ such that, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (A.2)
@@
Obviously, (A.2) holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $10<\gamma<$ 1. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $x,y\in\operatorname{int}\operatorname{dom}\phi$, $y\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
@@
Because $\phi$ is quadratic when $D_{\phi}$ is symmetric [Bauschke2001-ry, Lemma 3.16], it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
A.2 Lemmas of Smooth Adaptable Property

### v4

@@ 6 Numerical Experiments
In this section, we conduct numerical experiments to examine the performance of our algorithms. All numerical experiments were performed in Julia 1.11 using the FrankWolfe.jl package (Besançon et al., 2022) [^8] https://github.com/ZIB-IOL/FrankWolfe.jl on a MacBook Pro with an Apple M2 Max and 64GB LPDDR5 memory.
We compare BregFW (Algorithm 1) and BregAFW (Algorithm 3, whose results are shown in Figures 7) using the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, MD, and ProjGD (their notation is described in Appendix F). In this section, we consider $\ell_{p}$ loss problems and phase retrieval, and we report additional experiments on nonnegative linear inverse problems, low-rank minimization, and NMF in Appendix F. Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for the away-step FW algorithm with the open loop. Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see Braun et al. (2025). We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, $\gamma_{\max}=1$, and the termination criterion $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle\leq 10^{-7}$ throughout all experiments.
@@
Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$.
Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$.
@@ 6.1 $\ell_{p}$ Loss Problem
We consider the $\ell_{p}$ loss problem (Kyng et al., 2015; Maddison et al., 2021) to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Let $f(x)=\|Ax-b\|_{p}^{p}$. The gradient $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $1<p<2$, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi:=f$.
We consider real-world gas sensor data. [^9] The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$. We use the $\ell_{2}$ constraint $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|^{2}\leq b_{\max}\}$. We compare BregFW with EucFW, OpenFW, and ProjGD because ShortFW does not converge. The maximum number of iterations is 1000 for all methods except OpenFW, for which it is 5000. The initial point $x_{0}$ is an all-ones vector. We set $p=1.1$, i.e., $f$ is $\mathcal{C}^{1}$ not $\mathcal{C}^{2}$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $b_{\max}=130$ and $b_{\max}=200$, respectively. Because $\nabla f$ is not Lipschitz continuous, ShortFW and EucFW do not converge. The primal gap and the FW gap achieved by BregFW are the smallest among these methods because only BregFW has a theoretical guarantee for this problem.
@@ 6.2 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in P\subset\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval arises in many fields of science and engineering, such as image processing (Candés et al., 2015), X-ray crystallography (Patterson, 1934; 1944), and optics (Shechtman et al., 2015). To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i}(|\langle a_{i},x\rangle|^{2}-b_{i})^{2}$. Let $\phi$ be defined by $\sigma$-strongly convex $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ for $\sigma\leq 1$. For any $L$ satisfying $L\geq\sum_{i}(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ (Bolte et al., 2018, Lemma 5.1). In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. Thus, $L\geq\rho/\sigma$ holds with $\sigma=1$.
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compare BregFW with EucFW, ShortFW, and OpenFW. The maximum number of iterations is 1000 for all methods except OpenFW, for which it is 5000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized them to have norm 1 (random seed 1234). We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $(m,n)=(1000,10000)$ and $K=2000$ and for $(m,n)=(2000,10000)$ and $K=2000$, respectively. The primal gap and the FW gap by the adaptive Bregman step-size strategy are the smallest among these step-size strategies. In both cases, BregFW stopped before the 1000th iteration.
@@ 7 Conclusion and Future Work
We propose FW algorithms with an adaptive Bregman step-size strategy for smooth adaptable functions, which do not require Lipschitz continuous gradients. We have established convergence rates ranging from sublinear to linear, depending on various assumptions for convex and nonconvex optimization. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes and have established global convergence for convex optimization under the HEB condition and local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments show that our algorithms outperform existing algorithms.
Limitations and future work:
@@
Our step-size strategy for $L$-smad functions requires the value of $\nu$, which must be estimated because it depends on the choice of the Bregman distance. Theoretically, linear convergence of the FW algorithm always holds for convex optimization if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$. For $q>1+\nu$, if $t\leq t_{0}$, linear convergence holds; otherwise, sublinear convergence. The condition $q=1+\nu$, derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$. On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, i.e., the $2$-HEB condition. Since this is derived from the weak convexity of $f$, relaxing the quadratic growth condition may not be possible. One direction is to consider an extension using the difference of convex functions (DC) optimization. Recently, Maskan et al. (2025) incorporated DC optimization into FW algorithms and established sublinear convergence. Our algorithms could potentially benefit from DC optimization techniques.
Acknowledgments
@@
This project has been partially supported by the Japan Society for the Promotion of Science (JSPS) through Invitational Fellowships for Research in Japan program (L24503), JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351 and JST ERATO Grant Number JPMJER1903. Research reported in this paper was also partially supported by the Deutsche Forschungsgemeinschaft (DFG) through the DFG Cluster of Excellence MATH+ (grant number EXC-2046/1, project ID 390685689).
LLM Usage
We used large language models, ChatGPT-5/5.1/5.2, solely to polish the grammar and wording of the author-written text (e.g., to improve clarity and tone). The model did not generate novel ideas, analyses, code, or results. All outputs were reviewed and edited by the authors, who take full responsibility for the content.
Ethics Statement
This study uses only publicly available datasets that contain no personally identifiable information. We did not collect any new human-subject data; therefore, IRB approval was not required. We also reviewed the datasets for potentially harmful content.
Reproducibility Statement
We include all evaluation code, configuration files, and scripts in the supplementary material. We provide the exact preprocessing steps for all datasets used. We fix and disclose random seeds for all experiments. Results are reported as mean $\pm$ standard deviation over multiple runs.
Appendix A Related Work
Overview of Existing Bregman-Based Algorithm Studies:
The Bregman gradient method was first introduced as the mirror descent by Nemirovskij and Yudin (1983). Accelerated mirror descent methods were proposed by Krichene et al. (2015); Allen-Zhu and Orecchia (2017). Some first-order methods based on the Bregman distance do not require the global Lipschitz continuity of $\nabla f$. Bolte et al. (2018) proposed the Bregman proximal gradient algorithm (BPG) for nonconvex optimization problems and established its global convergence under the smooth adaptable property (see Definition 2.3). Hanzely et al. (2021) introduced the accelerated BPGs. Takahashi et al. (2022) developed the BPG for the DC optimization and its accelerated version. The subproblem of BPG is often not solvable in closed form. Takahashi and Takeda (2025); Fujiki et al. (2025) used approximations to Bregman distances to solve subproblems in closed form. Takahashi et al. (2026) developed the majorization-minimization BPG and its accelerated version, and applied them to NMF. Rebegoldi et al. (2018) proposed an inexact version of the Bregman proximal gradient algorithm.
Appendix B Results from Smooth Adaptable Property
B.1 Properties of Bregman Distances
@@
Example B.1.
Well-known choices for $\phi$ and $D_{\phi}$ are listed below; for more examples, see, e.g., (Bauschke et al., 2017, Example 1), (Bauschke and Borwein, 1997, Section 6), and (Dhillon and Tropp, 2008, Table 2.1).
@@
    The Boltzmann–Shannon entropy $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $0\log 0=0$ and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(x_{i}\log\frac{x_{i}}{y_{i}}-x_{i}+y_{i})$ is called the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951).
@@
    The Burg entropy $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}_{++}^{n}$. Then, $D_{\phi}(x,y)=\sum_{i=1}^{n}(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1)$ is called the Itakura–Saito divergence (Itakura and Saito, 1968).
@@
    Let $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. The Bregman distance $D_{\phi}$ is used in phase retrieval (Bolte et al., 2018), low-rank minimization (Dragomir et al., 2021), NMF (Mukkamala and Ochs, 2019), and blind deconvolution (Takahashi et al., 2023).
We recall the triangle scaling property for Bregman distances from (Hanzely et al., 2021, Section 2), where several properties of the triangle scaling property are also shown.
Definition B.2 (Triangle Scaling Property (Hanzely et al., 2021, Definition 2)).
@@
$D_{\phi}((1-\gamma)x+\gamma y,(1-\gamma)x+\gamma z)\leq\gamma^{\nu}D_{\phi}(y,z).$ (B.1)
Now, substituting $z\leftarrow x$ on the left-hand side of Equation B.1, we obtain
@@
We will now show that a stronger version can be obtained if $\phi$ is strictly convex, so that we can rephrase $\nu$ with $1+\nu$ where $\nu>0$, i.e., the right-hand side is superlinear, which will be crucial for the convergence analysis later.
Lemma B.3.
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, if $\phi$ is strictly convex, then there exists $\nu>0$ such that, for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and all $\gamma\in[0,1]$, it holds that
$D_{\phi}((1-\gamma)x+\gamma y,x)\leq\gamma^{1+\nu}D_{\phi}(y,x).$ (B.2)
@@
Equation B.2 holds if $y=x$ or $\gamma\in\{0,1\}$. In what follows, we assume $y\neq x$ and $0<\gamma<1$. Let $g(\nu):=\gamma^{1+\nu}D_{\phi}(y,x)-D_{\phi}((1-\gamma)x+\gamma y,x)$ for all $\in\operatorname{int}\operatorname{dom}\phi, yxy\neq x$ and all $\gamma\in(0,1)$. Using $D_{\phi}(y,x)>0$ due to the strict convexity of $\phi$, we have, for any $\nu\geq 0$,
@@
Because $\phi$ is quadratic when $D_{\phi}$ is symmetric (Bauschke and Borwein, 2001, Lemma 3.16), it always holds that $D_{\phi}((1-\gamma)x+\gamma y,x)=\gamma^{1+\nu}D_{\phi}(y,x)$ with $\nu=1$.
B.2 Properties of Smooth Adaptable Property